OpenStudy (anonymous):

square root of 5 ( square root of 15 + 3 square root of 2. a. 45 square root of 10 b. 9 square root of 10 c. 5 square root of 3 + 15 square root of 10 d. 5 square root of 3+ 3 square root of 10

OpenStudy (anonymous):

Is this the original equation? \[\sqrt{5}(\sqrt{15}+3\sqrt{2})\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

The distributive property says: \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

More precisely: \[\sqrt{a}(\sqrt{b} + c \sqrt{d}) = \sqrt{a \times b} + c \sqrt{a \times d}\]

OpenStudy (anonymous):

c ?

OpenStudy (anonymous):

Replace the letters in water's equation with the numbers in the original equation. What would that be? Don't worry about multiplying it out for now.

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

See the formula that i have written above you just want to substitute the value there.. a = 5, b = 15, c =3 and d=2..

OpenStudy (anonymous):

i got 35.467595040383.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

See, throw away your calculator.. There is no need of it.. tell me when you multiply root(5) and root(15) what will you get?

OpenStudy (anonymous):

75 ?

OpenStudy (anonymous):

75 in the root or simply 75?

OpenStudy (anonymous):

are you asking me what the root of 5 times the root of 15 equals or just 15 times 5 ?

OpenStudy (anonymous):

I am asking the first one that is root of 5 multiplied by root of 15?

OpenStudy (anonymous):

8.660254037844386

OpenStudy (anonymous):

You don't need a calculator for this whatsoever. Even the answer isn't an exact number. Punching it into a calculator will not help you.

OpenStudy (anonymous):

so how do i do this ? ughhh ! ): this is not helping me

OpenStudy (anonymous):